What is the force between two electrons 1.0 centimeter apart in air?

1
Expert's answer
2022-03-16T18:30:09-0400

"constant, k_e=9\times 10^9"

charge of electron "q_1=q_2=1.6\times 10^{-19}C"

distance, "d=1cm=0.01m"

Force, "F=\frac{k_e \times q_1 q_2}{d^2}=\frac{9\times 10^9 \times 1.6\times 10^{-19}\times1.6 \times 10^{-19}}{0.01^2}=2.304\times 10^{-24}N"